OpenStudy (anonymous):

What is the vertex form of the equation.. y=x^2-6x+8

OpenStudy (tkhunny):

You must "Complete the Square"! y = (x^2 - 6x + _____) + 8 - _____ With what shall we fill in BOTH blanks?

OpenStudy (anonymous):

9..

OpenStudy (tkhunny):

Yes!!! Excellent work. Put that in and simplify. Let's see what you get.

OpenStudy (anonymous):

y=(x^2-6+9)+8-9 y=x-3^2+8-9

OpenStudy (tkhunny):

Pretty close. Missing some parentheses and go ahead and add those last two values. \(y = (x-3)^{2} - 1\)